# QOSST: A Highly-Modular Open Source Platform for Experimental Continuous-Variable Quantum Key Distribution

### v1

@@ Acknowledgments
Figure 1 was created by partly using the gwoptics component library distributed under CC-BY-NC license, attribution to Alexander Franzen Alexander Franzen (2012). The authors acknowledge financial support from the European Unions’s Horizon Europe research and innovation programme under the project QSNP (Grant No. 101114043) and the PEPR integrated project QCommTestbed (ANR-22-PETQ-0011), which is part of Plan France 2030.
@@ Appendix A Calibration of Alice
A characterization of Alice’s VOA is necessary to know the relation between the applied voltage and the attenuation. In the experiment, a voltage of $2.9\;$\displaystyle\mathrm{V}\text{/}$$ is applied, corresponding to an attenuation of $3.78\;$\displaystyle\mathrm{dB}\text{/}$$. Both characterizations can be done using a script in the qosst-alice package.
@@ Appendix B Calibration of Bob
$\eta_{PD}=\frac{n_{e}}{n_{p}}=\frac{\frac{I}{e}}{\frac{P}{\frac{hc}{\lambda}}}=\frac{hc}{\lambda e}\mathcal{R}\stackrel{{\scriptstyle\lambda=1550\;$\displaystyle\mathrm{nm}\text{/}$}}{{\simeq}}\frac{\mathcal{R}}{1.25\;$\displaystyle\mathrm{A}\text{/}\mathrm{W}$}.$ (11)
For a balanced detector, the responsivity is usually defined as the sum of the two photocurrents divided by the input power at the beginning of the interferometer. For the whole detector, we can compute the responsivity as the sum of all photocurrents divided by the input optical power, and then divide this value by $1.25\;$\displaystyle\mathrm{A}\text{/}\mathrm{W}$$ to get the efficiency. In our setup, the Thorlabs balanced detector has two voltage monitoring outputs $V_{+},V_{-}$, such that
@@
where $I_{+}$ and $I_{-}$ are the photocurrents, $G_{m}=10\;$\displaystyle\mathrm{kV}\text{/}\mathrm{A}$$ is the monitoring gain and $V_{+}$ and $V_{-}$ are the voltages of the monitoring outputs. Hence
@@
The value of $\eta$ changes between the setup with the emulated channel using the VOA and the one with the fiber spool, since in the latter a manual polarisation controller is added to Bob’s setup. For both cases, $\eta$ is estimated using an automated script available in the qosst-bob package, giving $\eta_{voa}=55.4\%$ and $\eta_{fiber}=41.3\%$. We can compare those values with the expected efficiency of the receiver. The responsivity of the photodiodes in the balanced detector is typically around $0.9\;$\displaystyle\mathrm{A}\text{/}\mathrm{W}$$ at $1550\;$\displaystyle\mathrm{nm}\text{/}$$ (verified by measurement of the balanced detector only), giving an efficiency of $\eta_{PD}=72\%$. The excess losses are typically 7% (beam splitter), 15% (switch), 10% (polarisation maintaining mating sleeves, 2 in total), giving $\eta_{typ}=46\%$. Monitoring the photocurrents is crucial in order to get a good estimation of the detector efficiency $\eta$, which can lead to security issues if not properly estimated, in particular in the trusted detector model.

### v3

@@ Acknowledgments
Figure 1 was created by partly using the gwoptics component library distributed under CC-BY-NC license, attribution to Alexander Franzen [40]. The authors acknowledge financial support from the European Unions’s Horizon Europe research and innovation programme under the project QSNP (Grant No. 101114043) and the PEPR integrated project QCommTestbed (ANR-22-PETQ-0011), which is part of Plan France 2030.
@@ Appendix A Calibration of Alice
A characterization of Alice’s VOA is necessary to know the relation between the applied voltage and the attenuation. In the experiment, a voltage of $2.9\;$\mathrm{V}\text{/}$$ is applied, corresponding to an attenuation of $3.78\;$\mathrm{dB}\text{/}$$. Both characterizations can be done using a script in the qosst-alice package.
@@ Appendix B Calibration of Bob
$\eta_{PD}=\frac{n_{e}}{n_{p}}=\frac{\frac{I}{e}}{\frac{P}{\frac{hc}{\lambda}}}=\frac{hc}{\lambda e}\mathcal{R}\stackrel{{\scriptstyle\lambda=1550\;$\mathrm{nm}\text{/}$}}{{\simeq}}\frac{\mathcal{R}}{1.25\;$\mathrm{A}\text{/}\mathrm{W}$}.$ (11)
For a balanced detector, the responsivity is usually defined as the sum of the two photocurrents divided by the input power at the beginning of the interferometer. For the whole detector, we can compute the responsivity as the sum of all photocurrents divided by the input optical power, and then divide this value by $1.25\;$\mathrm{A}\text{/}\mathrm{W}$$ to get the efficiency. In our setup, the Thorlabs balanced detector has two voltage monitoring outputs $V_{+},V_{-}$, such that
@@
where $I_{+}$ and $I_{-}$ are the photocurrents, $G_{m}=10\;$\mathrm{kV}\text{/}\mathrm{A}$$ is the monitoring gain and $V_{+}$ and $V_{-}$ are the voltages of the monitoring outputs. Hence
@@
The value of $\eta$ changes between the setup with the emulated channel using the VOA and the one with the fiber spool, since in the latter a manual polarisation controller is added to Bob’s setup. For both cases, $\eta$ is estimated using an automated script available in the qosst-bob package, giving $\eta_{voa}=55.4\%$ and $\eta_{fiber}=41.3\%$. We can compare those values with the expected efficiency of the receiver. The responsivity of the photodiodes in the balanced detector is typically around $0.9\;$\mathrm{A}\text{/}\mathrm{W}$$ at $1550\;$\mathrm{nm}\text{/}$$ (verified by measurement of the balanced detector only), giving an efficiency of $\eta_{PD}=72\%$. The excess losses are typically 7% (beam splitter), 15% (switch), 10% (polarisation maintaining mating sleeves, 2 in total), giving $\eta_{typ}=46\%$. Monitoring the photocurrents is crucial in order to get a good estimation of the detector efficiency $\eta$, which can lead to security issues if not properly estimated, in particular in the trusted detector model.